ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 47 Zemánek, Jaroslav, Institute of Mathematics of the Polish Academy of Sciences, Poland Powers and resolvents We intend to discuss recent results and problems relating various properties of resolvents and powers of linear operators. Sample questions: (1) Find a (nonzero) quasinilpotent operator with resolvent majorized by const/|z| for Rez > 0. (2) Does sup_{n} n∥T^{n} − T^{n+1}∥ < ∞ imply sup_{n} ∥T^{n}∥ < ∞, where

n = 1, 2, …?